OpenStudy (cakt07):

Which statement is true? A. –16 + 23 = –17 + 10 B. 15 + (–5) > 23 + (–12) C. 8 + (–7) < –15 + 14 D. –12 + (–11) > 5 + (–29)

OpenStudy (cakt07):

@Mertsj

OpenStudy (gorv):

-12-11=-23 5-29=-24 -23>-24 D is true

OpenStudy (gorv):

@cakt07

OpenStudy (cakt07):

Thx ^^

OpenStudy (cakt07):

a few more?

OpenStudy (gorv):

wlcme :)

OpenStudy (cakt07):

What is the value of the variable that makes the statement true? c + (–8) + (–5) = 0 A. –13 B. –3 C. 3 D. 13

OpenStudy (gorv):

c-8-5=0 c-13=0 c=13

OpenStudy (cakt07):

I got B

OpenStudy (cakt07):

c- 8-5=0 -3 - 8 = 5 5-(-5)= 0

OpenStudy (cakt07):

OHHHH Nvm i got it now

OpenStudy (gorv):

okkkk